# Enhancing Security in Multi-Robot Systems through Co-Observation Planning, Reachability Analysis, and Network Flow

### v1

@@ II-G Line-segment-ellipse reachability constraint
For cases, where the tangent interpolation point ${p_{\mathcal{L}}^{\mathcal{E}}}$ stays within the segment (i.e. red region in Figure 2d, where ${p_{\mathcal{L}}^{\mathcal{E}}}$ lies between ${p_{1}^{\mathcal{E}}}$ and ${p_{2r}^{\mathcal{E}}}$), the constraint can be seen as a plane-ellipse constraint in Section II-F. Otherwise (i.e. brown region in Figure 2d, where ${p_{\mathcal{L}}^{\mathcal{E}}}$ lies outside ${p_{1}^{\mathcal{E}}}$ and ${p_{2b}^{\mathcal{E}}}$), the constraint can be seen as a point-ellipse constraint in Section II-E.
ADMM constraint 4 (Line-segment-ellipsoid reachability constraint)
[…]
$\mathcal{Z}$ =$=\{\mathbf{q}\in\mathbb{R}^{nm}:\lVert D(\mathbf{q})\rVert=0\},$ (46)
@@
where $D_{p_{1}}$ and $D_{p_{2}}$ are the point-ellipsoid constraint projection function with respect to ${p_{1}^{\mathcal{E}}}$ and ${p_{2}^{\mathcal{E}}}$ in (24), and $D_{p_{\mathcal{L}}^{\mathcal{E}}}$ is the plane-ellipsoid constraint with respect to frame ${\mathcal{L}^{\mathcal{E}}}$ in (36).
II-H Convex-polygon-ellipse reachability constraint
To keep an ellipse away from a convex polygon, first, we need to keep all segments of the hyperplanes outside the ellipse. Similar to (35), we define
ADMM constraint 5 (Convex-polygon-ellipsoid reachability constraint)
$\mathcal{Z}$ =$=\{\mathbf{q}\in\mathbb{R}^{nm}:\lVert D(\mathbf{q})\rVert=0\},$ (52)
$\Pi_{\mathcal{Z}}(\mathbf{z})$ $=0,$ (53)
@@ II-H Convex-polygon-ellipse reachability constraint
where $D_{seg}$ are the constraint functions for each line segment used to define convex polygon region. ADMM constraint 5 needs to be supplemented with a convex obstacle constraint for the polygon (as introduced in [6]) to prevent cases where the ellipse is a subset of the region.
II-I Secured planning results and limitation
We test the newly introduced security constraints to a map exploration problem, illustrated in Figure 3 in both simulations and on an experimental testbed, involving a 3-robot system assigned to explore a $10m\times 10m$ region with one obstacle, Zone 1, and two forbidden zones, Zone 2 and Zone 3. The maximum velocity constraint $v_{max}=0.5m/dt$, time horizon $T=20$.
@@ II-I Secured planning results and limitation
We employ the APMAPF solver ([2]) to generate a MAPF plan with a co-observation schedule in a grid-world application (Fig.3a). This result serves as the initial trajectory input for the ADMM solver. This obtained plan is then transformed to a continuous configuration space, addressing map exploration tasks. Co-observation schedules are set up using the APMAPF algorithm for two forbidden regions. Reachability constraints are added to ensure an empty intersection between all robots’ reachability regions during co-observations and the forbidden regions. It’s important to note that the secured attack-proof solution is initially infeasible in the continues setting, unless additional checkpoints (e.g., stationary security cameras) are incorporated (additional task added for agent 3 to visit the checkpoint at time $8$). Further solutions to avoid these checkpoints are discussed in Section III.
@@
The simulation result, shown in Fig. 3c, displays reachability regions as black ellipsoids, demonstrating empty intersections with Zones 2 and 3. Explicit constraints between reachability regions and obstacles are not activated, assuming basic obstacle avoidance capabilities in robots. The intersections between obstacles and ellipsoids, as observed between agent 3 and Zone 1, are deemed tolerable. All constraints are satisfied, and agents have effectively spread across the map for optimal exploration tasks.
(a) Grid-world trajectory.
(b) Continuous world trajectory
(c) Securd result
Figure 3: Trajectory design of an example map exploration task for a three-robot system with start locations and destinations fixed. Task are planned in a $8\times 8$ grid world. Zone 1 is obstacle, Zone 2 and Zone 3 are safe zones. Figure 3a shows the planned trajectory with co-observation schedule in grid-world. Figure 3b shows the unsecured trajectory design in continuous world optimized with respect to a map exploration task where Zone 1 and 2 are considered together as one obstacle. A potential path that could break the security by entering region 1 is highlighted in dashline. Figure 3c shows the secured trajectory with the incorporation of co-observation and reachability constraints.
II-I1 Limitations
@@
Our solution demonstrates the potential of planning with reachability and co-observation to enhance the security of multi-agent systems. However, two primary challenges need to be addressed. Firstly, achieving a co-observation and reachability-secured plan may not always be feasible when robots are separated by obstacles or forbidden regions, making it impossible for them to establish co-observation schedules or find a reachability-secured path. For example, Agent 3 in Figure 3c required an additional security checkpoint to create secured reachability areas. Secondly, meeting security requirements may impact overall system performance, as illustrated by the comparison between Figure 3b and 3c. The introduction of security constraints led to the top left corner being unexplored by any robots. This trade-off between security and system performance is particularly significant given that system performance plays a crucial role in the decision to employ multi-agent systems. These challenges are addressed in Chapter III to ensure the effective use of planning with reachability and co-observation in securing multi-agent systems.
@@ III Cross-trajectory co-observation planning
To address the feasibility and performance trade-offs, we propose to form sub-teams on each route, and setup additional co-observations both within the sub-team and across different sub-teams. These cross-trajectory co-observations allow sub-teams to obtain trajectories that are closer to the optimal (as shown in Figure 1), because they do not require the entire sub-team to meet with other teams for co-observations, thus providing more flexibility.
@@ III-A Problem overview
We start with an unsecured MRS trajectory with $n_{p}$ routes $\{q_{p}\}_{p=1}^{N_{p}}$ (the ADMM-based planner in Section II without security constraints is used, but other planners are applicable). Here, the state $q_{p}(t),p\in\{\mathcal{I}_{0},\ldots,\mathcal{I}_{p}\}$ represents the reference position of the $i$-th sub-team at time $t\in\{0,\dots,T\}$. Introducing redundant robots into the MRS, we assume a total of $n>n_{p}$ robots are available and organized into sub-teams through a time-varying partition $\mathcal{I}(t)=\cup_{p}\mathcal{I}_{p}(t)$ where robots in each sub-team $\mathcal{I}_{p}$ share the same nominal trajectory. To ensure the fulfillment of essential tasks, at least one robot is assigned to adhere to the reference trajectory. The remaining redundant $n_{p}-n$ robots are strategically utilized to enhance security. These robots focus on co-observations either within their respective sub-teams, or when necessary, deviate and join another sub-team to provide necessary co-observations. The objective of this problem is to formulate a strategy for this new co-observation strategy, named cross-trajectory co-observation that the resulting MRS plan meet the security in Remark 1 while minimizing the number of additional robots required.
We model the planned trajectory $\{q_{p}\}_{p=1}^{N_{p}}$ as a directed checkpoint graph $G{q}=(V_{q},E_{q})$, where vertices $V_{q}$ represent key locations requiring co-observation (as required by Remark 1), and edges $E_{q}=E_{t}\cup E_{c}$ connect pairs of vertices if possible paths exist. Inspired by [8], additional robots in the sub-team are treated as flows in the checkpoint graph. This transforms the co-observation planning problem into a network multi-flow problem, solvable with general Mixed-Integer Linear Programming techniques.
@@
This section presents the two components of our approach: constructing the checkpoint graph based on unsecured multi-robot trajectories, and the formulation and solution of the network multi-flow problem.
@@ III-B Rapidly-exploring Random Trees
To find edges between different nominal trajectories, we use the $*\texttt{RRT}^{\texttt{*}}$ [28] algorithm to find paths from a waypoint on one trajectory to multiple destination points (i.e., reference trajectories of other sub-teams, in our method). As a optimal path planning algorithm, $*\texttt{RRT}^{\texttt{*}}$ returns the shortest paths between an initial location and points in the free configuration space, organized as a tree. We assume that the generated paths can be travelled in both directions (this is used later in our analysis). Key functions from $*\texttt{RRT}^{\texttt{*}}$ that are also used during constructions of the checkpoint graph are
@@
This function assigns a non-negative cost (total travel distance in our application) to the unique path from the initial position to $v$
$\texttt{Parent}(v)$
This is a function that maps the vertex $v$ to $v^{\prime}\in V$ such that $(v^{\prime},v)\in E$.
Our objective here is to ascertain the existence of feasible paths instead of optimize specific tasks. While the ADMM based solver Section II presented earlier offers a broader range of constraint handling, RRT*’s efficient exploration of the solution space, coupled with its ability to incorporate obstacle constraints, makes it a fitting choice for building the checkpoint graph.
@@ III-C Checkpoint graph construction
In this section, we define and search for security checkpoints and how to use $*\texttt{RRT}^{\texttt{*}}$ to construct the checkpoint graph $G_{q}$.
III-C1 Checkpoints
Let $q_{p}\in nT$ represent the trajectory for sub-team $p$ with time horizon $T$, where $q_{p}(t)\in n$ denotes the reference waypoint at time $t$. We identify checkpoints within $q_{p}$ that require co-observations for securty, as outlined in Remark 1. Checkpoint $v_{pi}=(q_{pi},t_{i})$ signifies a required co-observation for sub-team $p$, defined by a location $q_{pi}=q_{p}(t_{i})$ and time $t_{i}$. For simplicity, let $\mathcal{I}_{v_{i}}$ denote the sub-team to which $v_{i}$ belongs, and $q_{v_{pi}}=q_{pi}$ and $t_{v_{pi}}$ be the corresponding waypoint and time for checkpoint $v_{pi}$. To ensure the security of the reference trajectory, the reachability region between consecutive checkpoints must avoid intersections with forbidden areas. This requirement can be formally stated as:
Remark 2
A set of checkpoints $V_{p}=\{v_{p0},\dots,v_{pT}\}$ (arranged in ascending order of $t_{v_{pi}}$) can secure the reference trajectory for sub-team $p$, if $\mathcal{E}(q_{v_{pi}},q_{v_{p(i+1)}},t_{v_{pi}},t_{v_{p(i+1)}})\cap F=\emptyset$ for every $i$, where $F$ is the union of all forbidden areas.
A heuristic approach is provided (Algorithm 1) to locate the checkpoints on given trajectories (an optimal solution would likely be NP-hard, while the approach below works well enough for our purpose).
Algorithm 1 Secure Checkpoint Generate for sub-team $p$
$v_{p0}\leftarrow(q_{p}(0),0)$; $v_{pT}\leftarrow(q_{p}(T),T)$
$V_{p}=\{v_{p0},v_{pT}\}$
$t_{0}\leftarrow 0$; $t_{1}\leftarrow T$
while $\mathcal{E}(q_{p}(t_{0}),q_{p}(t_{1}),t_{0},t_{1})\cap F\neq\emptyset$ or $t_{1}-t_{0}>1$ do
     $i\leftarrow 1$; $j\leftarrow 1$
     while $\mathcal{E}(q_{p}(t_{0}),q_{p}(t_{0}+i),t_{0},t_{0}+i)\cap F=\emptyset$ and $t_{0}+i>t_{1}$ do
         $v_{pf}\leftarrow(q_{p}(t_{0}+i),t_{0}+i)$
         $i\leftarrow i+1$
     end while
     $V_{p}\leftarrow V_{p}\cup v_{pf}$
     while $\mathcal{E}(q_{p}(t_{1}-j),q_{p}(t_{1}),t_{1}-j,t_{1})\cap F=\emptyset$ and $t_{0}+i\geq t_{1}-j$ do
         $v_{pb}\leftarrow(q_{p}(t_{1}-j),t_{1}-j)$
         $j\leftarrow j+1$
     end while
     if $t_{0}+i\neq t_{1}-j$ then
         $V_{p}\leftarrow V_{p}\cup v_{pb}$
     end if
end while
III-C2 Cross-trajectory edges
To enable co-observation cross different sub-team at checkpoints, we search for available connection paths (cross-trajectory edges) between checkpoints on different trajectories, allowing robots deviate from one sub-team to perform co-observation with a different sub-team.
Cross-trajectory edges $E_{c}=(v_{1},v_{2})$ define viable paths between two reference trajectories, where $\mathcal{I}_{v_{1}}\neq\mathcal{I}_{v_{2}}$ and at least one of $v_{1}$ and $v_{2}$ correspond to a security checkpoint $\cup_{p}V_{p}$. The cross-trajectory edges must also adhere to reachability constraints $\mathcal{E}(q_{v_{1}},q_{v_{2}},t_{v_{1}},t_{v_{2}})\cap F=\emptyset$ to ensure that no deviations into forbidden areas can occur during trajectory switches.

### v2

@@ II-G Convex-polygon-ellipse reachability constraint
When the tangent interpolation point $p_{\mathcal{L}}^{\mathcal{E}}$} stays within the segment (i.e. red region in Fig. 3d), the constraint is a plane-ellipse constraint in Section II-F. Otherwise (i.e. brown region in Fig. 3d), the constraint is a point-ellipse constraint in Section II-E.
ADMM constraint 5 (Line-segment-ellipsoid reachability constraint)
$D_{seg}(\mathbf{q})$ $=\begin{cases}D_{p_{1}}(\mathbf{q})&({p_{1}^{\mathcal{E}}}-{p_{2}^{\mathcal{E}}})^{\mathrm{T}}({p_{\mathcal{L}}^{\mathcal{E}}}-{p_{2}^{\mathcal{E}}})<0\\ D_{p_{2}}(\mathbf{q})&({p_{2}^{\mathcal{E}}}-{p_{1}^{\mathcal{E}}})^{\mathrm{T}}({p_{\mathcal{L}}^{\mathcal{E}}}-{p_{1}^{\mathcal{E}}})<0\\ D_{p_{\mathcal{L}}^{\mathcal{E}}}(\mathbf{q})&\textrm{otherwise}\end{cases}$ (34)
$\mathcal{Z}$ $=\{\mathbf{q}\in\mathbb{R}^{nm}:\lVert D_{seg}(\mathbf{q})\rVert=0\},$ (35)
@@
where $D_{p_{1}}$ and $D_{p_{2}}$ are the point-ellipsoid constraint projection function (11) with respect to $p_{1}^{\mathcal{E}}$} and $p_{2}^{\mathcal{E}}$}, and $D_{p_{\mathcal{L}}^{\mathcal{E}}}$ is the plane-ellipsoid constraint (22) with respect to frame $\mathcal{L}^{\mathcal{E}}$}. This constraint needs to be supplemented with a convex obstacle constraint for the polygon (i.e. keep foci waypoints outside the region, introduced in [5]) to prevent cases where the ellipse is a subset of the region.
II-H Secured planning results and limitation
@@
We employ the Attack-Proof MAPF (APMAPF) solver [2] on an 8 by 8 grid world with a similar setup to generate a MAPF plan with a co-observation schedule in a grid-world application (Fig. 4a). This result is transformed to a continuous configuration space and serves as the initial trajectory input for the ADMM solver with an additional task function targeting the map exploration task. Co-observation schedules are set up using the APMAPF algorithm for two forbidden regions. Reachability constraints are added to ensure an empty intersection between all robots’ reachability regions during co-observations and the forbidden regions. It is important to emphasize that an APMAPF solution does not guarantee existence, nor does it ensure a successful transition to the continuous configuration space. In this case, while an APMAPF solution may be found, the secured, attack-proof solution becomes infeasible in the continuous setting because the robots’ mobility is no longer restricted to adjacent grids. Additional security measures, such as stationary security cameras or surveillance robots, need to be incorporated. For instance, in this scenario, we deploy a security camera as an additional security measure to observe agent 3 at time $8$ to ensure security.
@@
The simulation result, shown in Fig. 4b, displays reachability regions as black ellipsoids, demonstrating empty intersections with region 2 and 3. Explicit constraints between reachability regions and obstacles are not activated, assuming basic obstacle avoidance capabilities in robots. The intersections between obstacles and ellipsoids, as observed between agent 3 and region 1, are deemed tolerable. All constraints are satisfied, and agents have effectively spread across the map for optimal exploration tasks.
(a) Grid world result
(b) Secured result
Figure 4: Trajectory design of a three-robot system based on the grid world result. Region 1 is an obstacle, region 2 and region 3 are forbidden regions. (4a) Result of the APMAPF algorithm in $8\times 8$ grid-world without the map exploring objective. (4b) The secured trajectory with the incorporation of co-observation generated by APMAPF and additional reachability constraints. Attack attempts into the forbidden region (red dashed line) will cause team 1 miss the next scheduled co-observation.
II-H1 Limitations
@@
Our solution demonstrates the potential of planning with reachability and co-observation to enhance the security of MRS. However, two primary challenges need to be addressed. Firstly, achieving a co-observation and reachability-secured plan is not always feasible, particularly when obstacles or restricted regions separate the robots, preventing them from establishing co-observation schedules or finding reachability-secured paths. For example, agent 3 in Fig. 4b requires additional security measures to create secured reachability areas. Secondly, security requirements can impact overall system performance, as illustrated by the comparison between Fig. 2b and Fig. 4b. The introduction of security constraints resulted in the top left corner remaining unexplored by any robots. This trade-off between security and system performance is particularly significant as system performance is a key factor in the decision of MRS deployments. These challenges are further addressed in Section III, ensuring the effective integration of reachability and co-observation in securing MRS.
@@ III Cross-trajectory co-observation planning
To address the feasibility and performance trade-offs, we propose to form sub-teams on each route, and setup additional co-observations both within the sub-team and across different sub-teams. These cross-trajectory co-observations allow sub-teams to obtain trajectories that are closer to the optimal (as shown in Fig. 1), because they do not require the entire sub-team to meet with other teams for co-observations, thus providing more flexibility.
@@ III-A Problem overview
We start with an unsecured MRS trajectory with $n_{p}$ routes $\{q_{p}\}_{p=1}^{N_{p}}$ (the ADMM-based planner in Section II without security constraints is used, but other planners are applicable). Here, the state $q_{p}(t),p\in\{\mathcal{I}_{0},\ldots,\mathcal{I}_{p}\}$ represents the reference position of the $p$-th sub-team at time $t\in\{0,\dots,T\}$. Introducing redundant robots into the MRS, we assume a total of $n>n_{p}$ robots are available and organized into sub-teams through a time-varying partition $\mathcal{I}(t)=\cup_{p}\mathcal{I}_{p}(t)$ where robots in each sub-team $\mathcal{I}_{p}$ share the same nominal trajectory. To ensure the fulfillment of essential tasks, at least one robot is assigned to adhere to the reference trajectory. The remaining redundant $n_{p}-n$ robots are strategically utilized to enhance security. These robots focus on co-observations either within their respective sub-teams, or when necessary, deviate and join another sub-team to provide necessary co-observations. The objective of this problem is to formulate a strategy for this new co-observation strategy, named cross-trajectory co-observation, such that the resulting MRS plan meets the security in Definition 4 while minimizing the number of additional robots required.
Our strategy is based on the following concept:
Definition 6
The $i$-th checkpoint $v_{pi}=(q_{p},t_{i})$ for sub-team $p$ is defined as a location $q_{pi}=q_{p}(t_{i})$ and time $t_{i}$. It represents either the first or last instance where a robot is co-observed by a teammate, and serves as a point where the robot can leave or join a new team.
For simplicity, let $\mathcal{I}_{v_{i}}$ denote the sub-team to which $v_{i}$ belongs. To ensure the security of the reference trajectory, the reachability region between consecutive checkpoints must avoid intersections with forbidden regions. This requirement can be formally stated as:
Remark 2
A set of checkpoints $V_{p}=\{v_{p0},\dots,v_{pT}\}$ (arranged in ascending order of $t_{v_{pi}}$) can secure the reference trajectory for sub-team $p$, if $\mathcal{E}(q_{v_{pi}},q_{v_{p(i+1)}},t_{v_{pi}},t_{v_{p(i+1)}})\cap F=\emptyset$ for every $i$, where $F$ is the union of all forbidden regions.
Definition 7
The planned trajectory $\{q_{p}\}_{p=1}^{N_{p}}$ is represented as a directed checkpoint graph $G=(V,E)$, where $V=(\cup_{p}V_{p})\cup V_{c}$ is the set of vertices representing waypoints on the planned trajectory and $E=E_{t}\cup E_{c}$ is the set of edges representing feasible paths connecting $V$. The components of the checkpoint graph are defined as follows:
Checkpoints $V_{p}$
where additional robots are required for security through co-observation.
Trajectory edges $E_{t}$
represent the original planned path of the primary robot.
Connection vertices $V_{c}$
represent non-checkpoint locations linked via $E_{c}$.
Cross-trajectory edges $E_{c}$
represent feasible paths where robots deviate from the original trajectory to join other teams or provide co-observation support, at least one endpoint be a checkpoint.
We assume that the planned trajectory for each sub-team has a robot operating on it. The co-observation planning problem is then transformed into a network multi-flow problem by modeling the additional robots as flows traveling through $G$. These flows either follow the planned trajectory via trajectory edges $E_{t}$ or switch teams via cross-trajectory edges $E_{c}$ to ensure valid and secured paths. Additionally, they are required to visit every checkpoint $V_{p}$ to meet the security requirements introduced in 2. This formulation reduces the co-observation problem to finding valid flow patterns in $G$. The problem is then formulated into linear objectives and constraints, enabling the use of Mixed-Integer Linear Programming (MILP) techniques to efficiently compute feasible solutions.
@@
This section presents the two components of our approach: constructing the checkpoint graph based on unsecured multi-robot trajectories and the formulation and solution of the network multi-flow problem.
@@ III-B Rapidly-exploring Random Trees
To find edges between different nominal trajectories, we use the $\texttt{RRT}^{\texttt{*}}$ [32] algorithm to find paths from a waypoint on one trajectory to multiple destination points (i.e., reference trajectories of other sub-teams, in our method). As an optimal path planning algorithm, $\texttt{RRT}^{\texttt{*}}$ returns the shortest paths between an initial location and points in the free configuration space, organized as a tree. We assume that the generated paths can be traveled in both directions (this is used later in our analysis). Key function from $\texttt{RRT}^{\texttt{*}}$ is used during the construction of the checkpoint graph, specifically
@@
This function returns a total travel distance to the unique path from the initial position to $v$.
Our objective here is to ascertain the existence of feasible paths instead of optimizing specific tasks. While the ADMM-based solver Section II presented earlier offers a broader range of constraint handling, RRT*’s efficient exploration of the solution space, coupled with its ability to incorporate obstacle constraints, makes it a fitting choice for building the checkpoint graph.
@@ III-C Checkpoint graph construction
In this section, we define and search for security checkpoints and how to use $\texttt{RRT}^{\texttt{*}}$ to construct the checkpoint graph $G_{q}$. We use a heuristic approach (Algorithm 1) to locate the checkpoints on given trajectories. Algorithm 1 is applied independently to each sub-team to generate their initial checkpoints.” While an optimal solution would likely be NP-hard, this approach works well enough for our purpose.
Algorithm 1 Checkpoint Graph Initialization for Team $\mathcal{I}_{p}$
1:Initialization: Set $t_{0}\leftarrow 0$ and $t_{3}\leftarrow T$, and add $(q_{p}(t_{0}),t_{0})$ and $(q_{p}(t_{3}),t_{3})$ to $V_{p}$ as start and end vertices.
2:while $\mathcal{E}(q_{p}(t_{0}),q_{p}(t_{3}),t_{0},t_{3})\cap F\neq\emptyset$ and $t_{0}\leq t_{3}$ do
3:  Forward Search: Find the largest $t_{1}\in\{t_{0}+1,\dots,t_{3}\}$ such that $\mathcal{E}(q_{p}(t_{0}),q_{p}(t_{1}),t_{0},t_{1})\cap F=\emptyset$. Once found, add $(q_{p}(t_{1}),t_{1})$ to $V_{p}$.
4:  Backward Search: Find the smallest $t_{2}\in\{t_{1},\dots,t_{3}-1\}$ such that $\mathcal{E}(q_{p}(t_{3}),q_{p}(t_{2}),t_{3},t_{2})\cap F=\emptyset.$ Once found, add $(q_{p}(t_{2}),t_{2})$ to $V_{p}$.
5:  Update Indices: Set $t_{0}\leftarrow t_{1}$ and $t_{3}\leftarrow t_{2}$.
6:end while
Figure 5: Example of Algorithm 1. Start at both ends, each trajectory finds the largest reachability region avoiding forbidden areas, then repeats the process on remaining segments to build secure checkpoints.
III-C1 Cross-trajectory edges
To enable co-observation across different sub-teams at checkpoints, we search for available connection paths (cross-trajectory edges) between checkpoints on different trajectories, allowing robots to deviate from one sub-team to perform co-observation with a different sub-team.
Cross-trajectory edges $E_{c}=(v_{1},v_{2})$ define viable paths between two reference trajectories, where $\mathcal{I}_{v_{1}}\neq\mathcal{I}_{v_{2}}$ and at least one of $v_{1}$ and $v_{2}$ corresponds to a security checkpoint $\cup_{p}V_{p}$. The cross-trajectory edges must also adhere to reachability constraints $\mathcal{E}(q_{v_{1}},q_{v_{2}},t_{v_{1}},t_{v_{2}})\cap F=\emptyset$ to ensure that no deviations into forbidden regions can occur while switching trajectories.